# Referential communication in heterogeneous communities of pre-trained visual deep networks

### v5

@@ 1 Introduction
After reviewing related work (Section 2) and presenting our general setup (Section 3), we delve into our experiments in Section 4. First, in Section 4.1, we show that it is indeed possible for sets of heterogeneous pre-trained networks to successfully converge on a referent through an induced communication protocol. In Section 4.2, we study referential generalization, showing that the developed protocol is sufficiently flexible that the networks can, to some extent, use it to refer to objects that were not seen during the training phase (the “unfamiliar mammal” case we discussed above), as well as to tell apart distinct instances of the same object (the “distinguishing between different dogs” case above). We further show that the networks can successfully communicate about images from new datasets that were not used during communication training. In Section 4.3, we consider instead agent generalization. We show that, given a community of networks that have been trained to communicate, it is faster for a new pre-trained network with a different architecture, training objective or training corpus to acquire the existing community protocol, than for the enlarged community to develop a new protocol from scratch. This suggests that we could use a small set of networks to develop a “universal” protocol, that can then be taught in a supervised way to new networks that require the envisaged communication functionality. In Section 4.4 we offer some qualitative and quantitative insights into the protocol, suggesting that the network populations have developed a single shared code, that is not (only) referring to low-level image features, but also to higher-level semantic concepts, which we identify using sparse autoencoders (SAEs). The Conclusion (Section 5) takes stock of what we found, and discusses the way ahead.
@@ 2 Related work
Representation similarity and model stitching
Our work touches upon the issue of whether different neural networks share similar representations (e.g., Li et al., 2016; Morcos et al., 2018; McNeely-White et al., 2020). Some of the work on this topic adopts a method known as “model stitching,” which consists in connecting intermediate layers of two models to measure their compatibility (Lenc & Vedaldi, 2019; Bansal et al., 2021). This can be seen a form of model-to-model communication. Moschella et al. (2023); Maiorca et al. (2023), in particular, have recently introduced an unsupervised method that allows heterogeneous encoder and decoder networks to be “stitched” together in a 0-shot manner, based on second-order object representations (objects are represented by their similarities to a set of reference objects). Applying this approach to our referential communication scenario is a possible direction for future work.
@@ 3.2.1 Pre-trained vision modules
As vision modules, we use widely used networks that have freely downloadable parameters (Table 1). We mainly consider architectural variations of networks trained with a supervised classification objective, as this is the most common axis of variation across modern models. In particular, we use both convolutional (ResNet: He et al. (2016), Inception: Szegedy et al. (2014), VGG: Simonyan & Zisserman (2015)) and attention-based networks (ViT: Dosovitskiy et al. (2021), Swin: Liu et al. (2021)), trained on ILSVRC2012 ImageNet data (Russakovsky et al., 2015). We however also include in our collection a ResNet (from Desai & Johnson (2021)) that was trained on a different dataset, namely COCO (Chen et al., 2015). The latter involves fewer labels as well as stronger data imbalance (Gauen et al., 2017). Finally, we include an attention-based network, DINO (Caron et al., 2021), that was trained on ImageNet, but using a self-supervised objective (denoted by its ViT-S/16 architecture in Table 1). Further pre-trained networks are used in the agent generalization experiments, as briefly discussed in Section 4.3 below: see Table 19 in Appendix K for details.
@@ 3.2.3 Training
Batch size is set at 64, the largest value we could robustly fit in GPU memory. As we sample distractors directly from training batches, on each training step the referential game is played 64 times, once with every different image in the batch as target and the other 63 images serving as distractors. All experiments are implemented using the EGG toolkit (Kharitonov et al., 2019). All parameters needed to reproduce our experiments with the toolkit (both the continuous setup reported here, and the discrete one discussed in Appendix C) can be found in Appendix E. Evidence of training convergence is provided in Appendix F in terms of gradient evolution profiles. Appendix G reports results for variants of training in which we used image augmentations and hard distractors.
@@ 3.3 Datasets
Agents are also tested on an out-of-domain (OOD) dataset containing classes from the larger ImageNet-21k repository, as pre-processed by Ridnik et al. (2021). We selected the new classes among those that are neither hypernyms nor hyponyms of ImageNet1k classes, and we made sure that no OOD class had a WordNet path similarity score [^1] As implemented in the NLTK toolkit: https://www.nltk.org/. above 0.125 with any ImageNet1k class. [^2] The COCO dataset, used to pre-train one of our visual modules, contains four labels slightly above the threshold, with maximum similarity at 0.17. We moreover used a number of heuristics, combined with manual inspection, to make sure the new classes were more or less at the same level of granularity as the ImageNet1k categories. After applying these filters, we were left with 52 classes denoting objects that belong to semantic domains disjoint from those represented in ImageNet1k, such as trees (olive tree, fir, red pine) and human professions (carpenter, physician, organ grinder). For each of them, we sampled all images available in the pre-processed ImageNet-21k dataset of Ridnik et al. (2021) (guaranteed to be at least 450). We used 90% of the OOD data for testing OOD communication accuracy (Section 4.2 below) and to train the classifiers in the experiments reported in Appendix H. The remaining 10% was used to test the latter classifiers. [^3] We provide scripts to reproduce our ImageNet1k and OOD datasets at ANONYMIZED-FOR-SUBMISSION.
@@ 4.1 Referential communication of homogeneous and heterogeneous networks
It is also possible to train the full agent population to very high accuracy in relatively few epochs (Population row in Table 2). In this setup, we use all 7 modules as both senders and receivers, randomly pairing them at training time, ensuring each agent interacts with every other agent. A common communication protocol emerges that is mutually intelligible across all 49 agent pairings. Note that the large increase in learning time compared to the one-to-one setups is only apparent, because during population training each agent is only seeing 1/7 of the data in each epoch. [^5] In other words, Table 2 is comparing on the same scale the time it takes to train 2 agents in the one-to-one setups to the time it takes to train 14 agents in the population setup. The difference is even sharper if we consider specific pairs of population agents, as at each turn an agent is paired with one of seven distinct partners. Thus, in terms of actual data exposure, a joint count of 28 (or 24 when using the 64-dimensional channel) epochs correspond to 4 (3.3) effective epochs of data exposure for a single agent, on average. Convergence is thus faster on average in population training than in one-to-one training.
@@ 4.2 Referential generalization
Having established with the previous experiment that it is possible to train a light communication layer that allows networks with heterogeneous architectures to successfully agree on a target referent, we go back to the questions about practical referent generalization we asked in the introduction. First, we explore whether the protocol can make more granular distinctions among objects coming from the same ImageNet1k classes that were used to develop it. For these purposes, we built a special test set where all distractors are from the same ImageNet1k class as the target (so that, say, one target magpie needs to be distinguished from other magpies, instead of swans and bullfrogs). We tested all 1,000 ImageNet1k classes in smaller batches of 32 candidates, to ensure we had enough images in all cases. In this single-class setup, randomly baseline accuracy is at around 3.1%. Note that, since we are interested in out-of-the-box zero-shot generalization, we use the very same protocols induced with random distractors in the experiments above, without any special adaptation to the single-class setup.
Table 3 shows that communication is at least partially successful at a more granular level than ImageNet1k classes, despite a large drop from the random-distractor performance reported above. There is a noticeable discrepancy between 16 and 64-dimensional messages, with the latter being generally better than the former, especially in the heterogeneous setup. We find this result intriguing, as we would have expected smaller message dimensionality to act as a regularizing bottleneck, leading to better generalization performance.
Table 3: Percentage accuracy on ImageNet Single-class and Out of Domain sets for two message dimensions.
Single-class (16) OOD (16) Single-class (64) OOD (64)
Homogeneous 47.2 $\pm$ 3.4 90.5 $\pm$ 8.0 47.4 $\pm$ 6.4 94.8 $\pm$ 10.4
Heterogeneous 31.0 $\pm$ 22.2 62.9$\pm$ 17.6 48.6 $\pm$ 2.1 82.2 $\pm$ 13.8
Population 34.7 $\pm$ 6.8 72.2 $\pm$ 13.8 36.5 $\pm$ 8.6 72.2 $\pm$ 44.8
@@
As Table 3 shows, there is a significant accuracy drop (larger in the more challenging heterogeneous and population setups), but performance remains good enough to suggest that the protocol is at least to some extent able to generalize to images depicting referents not seen during training (recall that chance performance is here at $1/64\approx{}1.6\%$). Larger dimensionality seems again to help generalization, at least in the homogeneous and heterogeneous setups.
Finally, we consider a different kind of referent generalization, in which we take the agents that were communication-trained on ImageNet, and test them on different datasets, that might contain very different referent distributions. In particular, Cifar100 focuses on a small number of concepts at a higher level of granularity than ImageNet; [^6] Note that, while the agents are trained on ImageNet1K data, Cifar100 was used for the limited hyperparameter searches we describe in Appendix D. Places205 focuses on scenes rather than objects; CelebA is entirely composed of celebrity faces.
Table 18 shows that 64-dimensional communication is still possible in this zero-shot dataset-transfer experiment. With chance performance still at $1/64\approx{}1.6\%$, on Places205 we observe even higher accuracy than on the ImageNet OOD set. As can be seen in examples of Fig. 3 and in Appendix N, Fig. 17, this might be at least partially explained by the presence of familiar objects in the scenes depicted in this dataset. Performance on the CelebA dataset is near-perfect for homogeneous agents, and comparable to that obtained on the single-class set for the other agents. This makes sense, as this face-only dataset can be seen as an extreme case of the single-class setup. We conjecture that heterogeneous population pairs are less likely to converge on low-level features that might help catching the more granular differences of the single-class datasets. Finally, it is surprising that, while still much higher than random, performance is lowest on Cifar100, which is, like ImageNet, an object-centric dataset. By inspecting the Cifar100 images, we realized that they have a much lower resolution than those of ImageNet, looking much blurrier. As we will see below in Section 4.4.3, the communication protocol is not robust to image blurring.
Results for 16-dimensional communication are available in Appendix J. They follow a similar pattern but with lower accuracies, especially in the homogeneous pair setup. We observe that the extra dimensions have a low impact on in-domain performance (+ 2% for all three agent pairing setups), but are very useful for referent generalization in all cases, reaching a maximum of + 75% for homogeneous pairs on Places205.
@@ 4.3 Agent generalization
To test ease of learning of a communication protocol, we add a single network, which we call the learner agent, to an existing community, either as sender or as receiver. [^7] Alternatively, a single pair of agents could initially be trained, and then all other agents could be sequentially added to the growing community. As we obtain comparable results when adding each learner in turn to all possible network pairs or sextuples, we reasonably expect that this alternative setup would also lead to similar outcomes. This new agent has to play the referential game with the original agents and learn their protocol. To single out ease of learning of the communication protocol itself, and not the original agents’ adaptability, we freeze all communication layer parameters except for the learner agent’s. The new communication layer is therefore the only one being updated. The focus of this experiment is on the heterogeneous learners, where the new agent’s vision module is different from that of the original agents. We nonetheless observe that, in the homogeneous case (not reported here), communication protocols are learned very quickly, needing at most one epoch to reach perfect or near-perfect accuracy.
@@
Test accuracy and learning speed of the learner agents are reported in Fig. 2. In all runs with a 64-dimensional communication channel, the learner agent succeeds at learning the communication protocol with accuracy of at least 90%. There is some more variance with 16-dimensional messages, but in this case as well nearly all runs converge to very high accuracy. Receivers learn better and faster on average (94%/95% for one-to-one with 16/64-messages, respectively, and 96%/97% for population) than senders (88%/95%and 95%/97%, respectively).
@@
We further replicate the agent generalization experiments with other pre-trained models that had not been used at any point during our experiments so far: LeViT: Graham et al. (2021), CaiT: Touvron et al. (2021), ResNeXt: Xie et al. (2017), DenseNet: Huang et al. (2017). Accuracies are above 95% from the first epoch for all settings (see Appendix K), strongly confirming the generality of the protocol, and suggesting that, once the latter has been developed, it can easily be transferred to novel architectures.
We also experiment with a visual module that has been trained with a natural-language-mapping objective (CLIP Radford et al. (2021)), since this might have introduced biases that makes it very different from purely visual models. We find that this is not the case. No drop in performance is observed for generalization of CLIP-induced protocols to new agents, or to new data: adaptation to this potentially outlying representation is not slower or less efficient (see Appendix L).

### v7

@@ 1 Introduction
After reviewing related work (Section 2) and presenting our general setup (Section 3), we delve into our experiments in Section 4. First, in Section 4.1, we show that it is indeed possible for sets of heterogeneous pre-trained networks to successfully converge on a referent through an induced communication protocol. In Section 4.2, we study referential generalization, showing that the developed protocol is sufficiently flexible that the networks can use it to refer to objects that were not seen during the training phase (the “unfamiliar mammal” case we discussed above), as well as to tell apart distinct instances of the same object (the “distinguishing between different dogs” case above). We further show that the networks can successfully communicate about images from new datasets that were not used during communication training. In Section 4.3, we consider instead agent generalization. We show that, given a community of networks that have been trained to communicate, it is faster for a new pre-trained network with a different architecture, training objective or training corpus to acquire the existing community protocol, than for the enlarged community to develop a new protocol from scratch. This suggests that we could use a small set of networks to develop a “universal” protocol, that can then be taught in a supervised way to new networks that require the envisaged communication functionality. In Section 4.4 we offer some qualitative and quantitative insights into the protocol, suggesting that the network populations have developed a single shared code, that is not (only) referring to low-level image features, but also to higher-level semantic concepts, which we identify using sparse autoencoders (SAEs). The Conclusion (Section 5) takes stock of what we found, and discusses the way ahead.
The main contributions of this paper are as follows:
  • We introduce and motivate the task of referential communication between pre-trained heterogeneous visual networks;
  • We show that a low-dimensional communication layer trained in a self-supervised way is sufficient to achieve high performance on the task, given a variety of different networks tested on in-domain data;
  • We show that the learned communication protocol partially transfers to out-of-domain data, including new and more granular referents than those seen at training time;
  • We show that a new pre-trained network can successfully and efficiently learn an existing protocol;
  • Quantitative and qualitative analysis of the protocol suggests that it is encoding high-level semantic features.
@@ 2 Related work
Representation similarity, model stitching and multimodal representation learning
Our work touches upon the issue of whether different neural networks share similar representations (e.g., Li et al., 2016; Morcos et al., 2018; McNeely-White et al., 2020; Huh et al., 2024). Some of the work on this topic adopts a method known as “model stitching,” which consists in connecting intermediate layers of two models to measure their compatibility (Lenc & Vedaldi, 2019; Bansal et al., 2021). This can be seen a form of model-to-model communication. Moschella et al. (2023); Maiorca et al. (2023), in particular, have recently introduced an unsupervised method that allows heterogeneous encoder and decoder networks to be “stitched” together in a 0-shot manner, based on second-order object representations (objects are represented by their similarities to a set of reference objects). Applying this approach to our referential communication scenario is a possible direction for future work. Research on multimodal representation learning and alignment (e.g., Darvish et al., 2023; Girdhar et al., 2023; Liang et al., 2023; Yu et al., 2024) is also strongly related to our aims. For example, Moayeri et al. (2023) present a cross-visual-model alignment method similar to our communication layer. However, differently from their research line, where the emphasis is on multiple modalities and to train a single system that works across them, we are specifically interested in connecting heterogeneous pre-trained visual models in a light and efficient way, and in studying the properties of the emergent protocol shared by a whole population of such models.
@@ 3.2.1 Pre-trained vision modules
As vision modules, we use widely adopted networks that have freely downloadable parameters (Table 1). We mainly consider architectural variations of networks trained with a supervised classification objective, as this is the most common axis of variation across modern models. In particular, we use both convolutional (ResNet: He et al. (2016), Inception: Szegedy et al. (2014), VGG: Simonyan & Zisserman (2015)) and attention-based networks (ViT: Dosovitskiy et al. (2021), Swin: Liu et al. (2021)), trained on ILSVRC2012 ImageNet data (Russakovsky et al., 2015). We however also include in our collection a ResNet (from Desai & Johnson (2021)) that was trained on a different dataset, namely COCO (Chen et al., 2015). The latter involves fewer labels as well as stronger data imbalance (Gauen et al., 2017). Finally, we include an attention-based network, DINO (Caron et al., 2021), that was trained on ImageNet, but using a self-supervised objective (denoted by its ViT-S/16 architecture in Table 1). Further pre-trained networks are used in the agent generalization experiments, as briefly discussed in Section 4.3 below: see Table 22 in Appendix M for details.
@@ 3.2.3 Training
Batch size is set at 64, the largest value we could robustly fit in GPU memory. As we sample distractors directly from training batches, on each training step the referential game is played 64 times, once with every different image in the batch as target and the other 63 images serving as distractors. All experiments are implemented using the EGG toolkit (Kharitonov et al., 2019) [^1] https://github.com/facebookresearch/EGG, scripts from our experiments are in ./egg/zoo/pop. All parameters needed to reproduce our experiments with the toolkit (both the continuous setup reported here, and the discrete one discussed in Appendix C) can be found in Appendix E. Evidence of training convergence is provided in Appendix F in terms of gradient evolution profiles. Appendix G reports results for variants of training in which we used image augmentations and hard distractors.
@@ 3.3 Datasets
Agents are also tested on an out-of-domain (OOD) dataset containing classes from the larger ImageNet-21k repository, as pre-processed by Ridnik et al. (2021). We selected the new classes among those that are neither hypernyms nor hyponyms of ImageNet1k classes, and we made sure that no OOD class had a WordNet path similarity score [^2] As implemented in the NLTK toolkit: https://www.nltk.org/. above 0.125 with any ImageNet1k class. [^3] The COCO dataset, used to pre-train one of our visual modules, contains four labels slightly above the threshold, with maximum similarity at 0.17. We moreover used a number of heuristics, combined with manual inspection, to make sure the new classes were more or less at the same level of granularity as the ImageNet1k categories. After applying these filters, we were left with 52 classes denoting objects that belong to semantic domains disjoint from those represented in ImageNet1k, such as trees (olive tree, fir, red pine) and human professions (carpenter, physician, organ grinder). For each of them, we sampled all images available in the pre-processed ImageNet-21k dataset of Ridnik et al. (2021) (guaranteed to be at least 450). We used 90% of the OOD data for testing OOD communication accuracy (Section 4.2 below) and to train the classifiers in the experiments reported in Appendix H. The remaining 10% was used to test the latter classifiers. [^4] We provide scripts to reproduce our ImageNet1k and OOD datasets at https://github.com/mahautm/emecom_pop_data
@@ 4.1 Referential communication of homogeneous and heterogeneous networks
It is also possible to train the full agent population to very high accuracy in relatively few epochs (Population row in Table 2). In this setup, we use all 7 modules as both senders and receivers, randomly pairing them at training time, ensuring each agent interacts with every other agent. A common communication protocol emerges that is mutually intelligible across all 49 agent pairings. Note that the large increase in learning time compared to the one-to-one setups is only apparent, because during population training each agent is only seeing 1/7 of the data in each epoch. [^6] In other words, Table 2 is comparing on the same scale the time it takes to train 2 agents in the one-to-one setups to the time it takes to train 14 agents in the population setup. The difference is even sharper if we consider specific pairs of population agents, as at each turn an agent is paired with one of seven distinct partners. In Appendix J, we show that, if we train the population for 7 times more epochs, so that each population agent sees as much data as an agent in the one-to-one settings, population accuracy further slowly increases, although it still generally lags behind that of one-to-one training. Thus, in terms of actual data exposure, a joint count of 28 (or 24 when using the 64-dimensional channel) epochs correspond to 4 (3.3) effective epochs of data exposure for a single agent, on average. Convergence is thus faster on average in population training than in one-to-one training.
@@ 4.2 Referential generalization
Having established with the previous experiment that it is possible to train a light communication layer that allows networks with heterogeneous architectures to successfully agree on a target referent, we go back to the questions about referent generalization we asked in the introduction. First, we explore whether the protocol can make more granular distinctions among objects coming from the same ImageNet1k classes that were used to develop it. For these purposes, we built a special test set where all distractors are from the same ImageNet1k class as the target (so that, say, one target magpie needs to be distinguished from other magpies, instead of swans and bullfrogs). We tested all 1,000 ImageNet1k classes in smaller batches of 32 candidates, to ensure we had enough images in all cases. In this single-class setup, random baseline accuracy is at around 3.1%. Note that, since we are interested in out-of-the-box zero-shot generalization, we use the very same protocols induced with random distractors in the experiments above, without any special adaptation to the single-class setup.
Tables 3 and 4 show that communication is at least partially successful at a more granular level than ImageNet1k classes, despite a large drop from the random-distractor performance reported above. There is a noticeable discrepancy between 16 (Table 3) and 64-dimensional messages (Table 4), with the latter being generally better than the former, especially in the heterogeneous setup. We find this result intriguing, as we would have expected smaller message dimensionality to act as a regularizing bottleneck, leading to better generalization performance.
16-dimensional Single-class OOD
Homogeneous 47.2 $\pm$ 3.4 94.8 $\pm$ 10.4
Heterogeneous 31.0 $\pm$ 22.2 82.2 $\pm$ 13.8
Population 34.7 $\pm$ 6.8 72.2 $\pm$ 44.8
Table 3: Percentage accuracy on ImageNet Out of Domain sets for 16-dimensional messages.
64-dimensional Single-class OOD
Homogeneous 47.4 $\pm$ 6.4 90.5 $\pm$ 8.0
Heterogeneous 48.6 $\pm$ 2.1 62.9$\pm$ 17.6
Population 36.5 $\pm$ 8.6 72.2 $\pm$ 13.8
Table 4: Percentage accuracy on ImageNet Single-class and Out of Domain sets for 64-dimensional messages
@@
As Tables 3 and 4 show, compared to in-domain results (Table 2), there is a significant accuracy drop (larger in the more challenging heterogeneous and population setups), but performance remains good enough to suggest that the protocol is able to generalize to images depicting referents not seen during training well above chance level (recall that chance performance is here at $1/64\approx{}1.6\%$). Larger dimensionality seems again to help generalization, at least in the homogeneous and heterogeneous setups.
Finally, we consider a different kind of referent generalization, in which we take the agents that were communication-trained on ImageNet, and test them on other datasets, that might contain very different referent distributions. In particular, Cifar100 focuses on a small number of concepts at a higher level of granularity than ImageNet; [^7] Note that, while the agents are trained on ImageNet1K data, Cifar100 was used for the limited hyperparameter searches we describe in Appendix D. Places205 focuses on scenes rather than objects; CelebA is entirely composed of celebrity faces.
Table 5 shows that 64-dimensional communication is still possible in this zero-shot dataset-transfer experiment (see Appendix K for the 16-dimensional case). With chance performance still at $1/64\approx{}1.6\%$, with Places205 we observe even higher accuracy than with the ImageNet OOD set. As shown in the examples of Fig. 3 and in Appendix P, Fig. 16, this might be at least partially explained by the presence of familiar objects in the scenes depicted in this dataset. Performance on the CelebA dataset is near-perfect for homogeneous agents, and comparable to that obtained on the single-class set for the other agents. This makes sense, as this face-only dataset can be seen as an extreme case of the single-class setup. We conjecture that heterogeneous population pairs are less likely to converge on low-level features that might help catching the more granular differences of the single-class datasets. Finally, it is surprising that, while still much higher than random, performance is lowest on Cifar100, which is, like ImageNet, an object-centric dataset. By inspecting the Cifar100 images, we realized that they have a much lower resolution than those of ImageNet, looking much blurrier. As we will see below in Section 4.4.3, the communication protocol is not robust to image blurring.
Results for 16-dimensional communication are presented in Appendix K. They follow a similar pattern but with lower accuracies, especially in the homogeneous pair setup. We observe that the extra dimensions have a low impact on in-domain performance (+ 2% for all three agent pairing setups), but are very useful for referent generalization in all cases, reaching a maximum of + 75% for homogeneous pairs in Places205.
Overall, the main takeaways from the referent generalization experiments are the following. First, out-of-domain communication works out of the box with results way above chance, although there is a drop with respect to in-domain communication. Second, generalization is more successful when generalizing to new objects and scenes than when the agents must distinguish between objects of the same class
@@ 4.3 Agent generalization
To test ease of learning of a communication protocol, we add a single network, which we call the learner agent, to an existing community, either as sender or as receiver. [^8] Alternatively, a single pair of agents could initially be trained, and then all other agents could be sequentially added to the growing community. Appendix L shows that the all-at-once approach we report here outperforms sequential population training. This new agent has to play the referential game with the original agents and learn their protocol. To single out ease of learning of the communication protocol itself, and not the original agents’ adaptability, we freeze all communication layer parameters except for the learner agent’s. The new communication layer is therefore the only one being updated. The focus of this experiment is on the heterogeneous learners, where the new agent’s vision module is different from that of the original agents. We nonetheless observe that, in the homogeneous case (not reported here), communication protocols are learned very quickly, needing at most one epoch to reach perfect or near-perfect accuracy.
@@
Test accuracy and learning speed of the learner agents are reported in Fig. 2. In all runs with a 64-dimensional communication channel, the learner agent succeeds at learning the communication protocol with accuracy of at least 90%. There is some more variance with 16-dimensional messages, but in this case as well nearly all runs converge to very high accuracy. Receivers learn better and faster on average (94%/95% for one-to-one with 16/64-messages, respectively, and 96%/97% for population) than senders (88%/95% and 95%/97%, respectively). We don’t have an explanation for this, but we believe it is related to the fact that senders must perform more exploration when associating the target image to a message in the emergent protocol, whereas the role of the receivers is more deterministic, in that they must just decode the sender messages (see Rita et al., 2020, for sender/receiver learning asymmetries in emergent communication).
@@
We further replicate the agent generalization experiments with other pre-trained models that had not been used at any point during our experiments so far: LeViT: Graham et al. (2021), CaiT: Touvron et al. (2021), ResNeXt: Xie et al. (2017), DenseNet: Huang et al. (2017). Accuracies are above 95% from the first epoch for all settings (see Appendix M), strongly confirming the generality of the protocol, and suggesting that, once the latter has been developed, it can easily be transferred to novel architectures.
We also experiment with a visual module that has been trained with a natural-language-mapping objective (CLIP Radford et al. (2021)), since this might have introduced biases that makes it very different from purely visual models. We find that this is not the case. No drop in performance is observed for generalization of CLIP-induced protocols to new agents, or to new data: adaptation to this potentially outlying representation is not slower or less efficient (see Appendix N).